You purchase one Microsoft July $72 put contract for a premium of $1.32. What is your maximum possible profit?
Answer #1

The put option is profitable if the stock price reduces below the strike price

Strike Price = X = $72

Premium Paid = P = $1.32

Let Stock Price = S

Profit = (X - S) - P

The profit is maximized when Stock Price S = 0

=> Max Profit = (72 - 0) - 1.32 = $70.68

Hence, profit per lot = 70.68*100 = $7,068